What is the answer to this equation?
X+Y=14
Y=X+4

Answer by jorel1380(3719)

X +Y=14
Y=X+4
Subsituting for Y, we get:
X+(X+4)=14
2X+4=14
2X=10
X=5
Y=X+4=9.